OpenStudy (anonymous):

Can Someone Help Me Check My Work...I Did it all i just want to know if its right :D Check Comment Box :D

OpenStudy (anonymous):

1st Solve by using the perfect squares method. x^2 + 8x + 16 = 0 2 -2 4 -4

OpenStudy (anonymous):

My answer was -4

OpenStudy (anonymous):

You're correct.

OpenStudy (anonymous):

Cool thank you :D

OpenStudy (anonymous):

Solve. x^2 + 5x + 6 = 0 (Points : 4) {2, 3} {-2, -3} {1,- 6} {-1, 6}<---- My answer

OpenStudy (phi):

x^2 + 5x + 6 = 0 the 2nd + sign (in front of the 6) means the two roots have the SAME SIGN (- or +) the + in front of the 5x means the larger factor is + you could check your answer by substituting it into the equation. Does -1 work? (I always start with the easiest number) -1*-1 + 5*-1 +6 2 -5 +6 3 not zero, so -1 is not a root. to find the factors, list out the factors of 6 (the last number in the equation) 1, 6 2, 3 notice that 2+3=5, (matches the 5 in 5x) so (x+2)(x+3) are the factors and the roots are x= -2 and x= -3 Let's check the -2: -2*-2 + 5*-2 +6 = 4-10+6= -6+6=0 Yes, it works. we could check x= -3 if we really wanted to be sure!